OpenStudy (anonymous):

-3/4 X n = -34 What is the value of n ?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \frac{\LARGE-3 }{\LARGE 4} \times n=-34 }\) . Like this?

OpenStudy (anonymous):

negative 3/4

OpenStudy (solomonzelman):

that is what I wrote, no?

OpenStudy (anonymous):

yeah same thing

OpenStudy (solomonzelman):

I will multiply both sides times: \(\large\color{slate}{ \color{red}{ \frac{\LARGE4 }{\LARGE -3}} }\) Like this, \(\large\color{slate}{ \frac{\LARGE-3 }{\LARGE 4} \color{red}{\times \frac{\LARGE4 }{\LARGE -3}}\times n=-34\color{red}{\times \frac{\LARGE4 }{\LARGE -3}} }\)

OpenStudy (solomonzelman):

Simplify each side for me please.

OpenStudy (anonymous):

one sec i need paper and pen

OpenStudy (anonymous):

n = -34 x -4/3

OpenStudy (anonymous):

4/-3

OpenStudy (solomonzelman):

you got right that the left side will have just n. but for the right side, what do you get?

OpenStudy (anonymous):

-34/1 x 4/-3

OpenStudy (solomonzelman):

yes very good for the first step.

OpenStudy (solomonzelman):

then multiply the tops and the bottoms

OpenStudy (solomonzelman):

-34 \(\large\color{slate}{\times }\) (-4/3) >>> (-34 \(\large\color{slate}{\times }\) -4) /3

OpenStudy (solomonzelman):

so your n will be equivalent to?

OpenStudy (anonymous):

ok 136/3

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

you're done with this Q. yw!

OpenStudy (anonymous):

n = 136/3

OpenStudy (anonymous):

thx